# "I’m not for sale" – Perceptions and limited awareness of privacy risks by digital natives about location data

### v1

@@ 3.5 Data analysis
The analysis was performed by two independent annotators, and a consensus was reached at the end of the process. For most fields, a full thematic analysis was neither required nor necessarily desirable; curation and normalisation of the data was however necessary. Indeed, stopping at the third step was enough for several fields, such as the reasons given to revoke location permission. A full thematic analysis was performed on the answers to Q32 “How do you think location data could be misused by a malicious person?”, the themes of which were loosely derived from (Citron and Solove 2021) (Physical harms, Economic harms, Reputational harms, Informational integrity, Psychological harms), see Table 1 in Appendix C.
We regrouped our codes for Q36 “What are the measures imposed by the GDPR that reduce these risks?” under general legal categories (Rights, Consent, Clear purpose and transparency, Misc.) for statistical purposes, see Table 2 in Appendix C. The rationale behind Q36 was to evaluate participants’ knowledge about the GDPR.
3.6 General Statistics
@@
Note that our population is within average with respect to their privacy concerns ($\mu$ = 4.2, $\sigma$ = 1.05) [^12] See (Zufferey et al. 2023), for an example of less concerned participants ($\mu$ = 3.5, $\sigma$ = 1.6) and (Naeini et al. 2017) for more concerned participants with $\mu$ between 4.79 and 5.44 brought down to a [0-6] scale. in spite of young adults being generally less concerned about their privacy than older populations (as mentioned in Section 2).
Figure 4: Plot of the Internet User’s Information Privacy Concerns (IUIPC) distribution of participants.
4 Results
@@
4.1 Privacy risky practices
@@
Figure 5: Participants tend to underestimate the number of apps that have access to the location (Q3 against Q8).
@@
Figure 6: Participants tend not to think about disabling location access for apps that are not actively used or forget to do so (Q11).
@@
4.2 Participants are poorly aware of the risks
@@
Figure 7: Proportion of participants aware of source of data from which location can be inferred (Q12).
@@
Figure 8: Answers to Q13 “Do you think it is possible to infer the following information from the location?”.
@@
4.3 The risk demonstrator increased awareness of participants about privacy
@@
8
Figure 9: The most worrisome inferences are informational harms (i.e., surveillance, stalking, inference, data selling, targeted advertising) (Q32).
@@ 5.1 Recommendations
5.2 Limitations
First, our demographics do not represent a fair sample of digital users. Our participants were mostly males, all of them young adults educated in a top European engineering school, although with different backgrounds and technical skills. This bias in our representativeness played a part in the socio-economic backgrounds of participants, who are most likely to come from well-off social environments. Recruiting participants in universities often faces this limitation in academic studies on usages. For instance, (Kulyk et al. 2022) involved 40 participants mostly young and well-educated, while (Karegar, Pettersson, and Fischer-Hübner 2020) involved 80 participants, mostly young, with 45% in the 18–25 age group and 20% between the ages of 26 and 35, a majority ($n=54$) were bachelor students or had their bachelor degrees. (Almuhimedi et al. 2015) involved 23 participants (65% female; ages 18–44, median=23), and (Kaushik et al. 2021) involved 392 participants ranging from 25 to 35.
@@ 5.2 Limitations
Second, the scope of our work must be restricted to a specific type of demonstrator, our results should not be generalized to all visualization and awareness tools. Moreover, the results largely presented self-reported and planned behaviors. Therefore, an interesting research avenue for future work is to study the actual behavioral changes introduced by the experiment of the tool, and long-term behavior analysis. Perceptions can also change according to the type of apps and its access to location data, as well as other sensitive personal data collected by mobile apps.
Third, participants used example data on the demonstration platform, which is perhaps less illustrative than if they had inspected their own location data. We envision working with personalized examples in the future. Our results also lack perspective on whether our participants were able to deeply understand the goal and the meaning of obfuscated traces and the epsilon number (see (Karegar, Alaqra, and Fischer-Hübner 2022)).
6 Conclusion
@@
Appendix A Paper Checklist to be included in your paper
  1. 1.
    For most authors…
    1. (a)
      Would answering this research question advance science without violating social contracts, such as violating privacy norms, perpetuating unfair profiling, exacerbating the socio-economic divide, or implying disrespect to societies or cultures? Yes, see Section 3.2 on Ethical Considerations
    2. (b)
      Do your main claims in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    3. (c)
      Do you clarify how the proposed methodological approach is appropriate for the claims made? Yes
    4. (d)
      Do you clarify what are possible artifacts in the data used, given population-specific distributions? Yes, see Section 3.5
    5. (e)
      Did you describe the limitations of your work? Yes, in Section 5.2
    6. (f)
      Did you discuss any potential negative societal impacts of your work? NA
    7. (g)
      Did you discuss any potential misuse of your work? NA
    8. (h)
      Did you describe steps taken to prevent or mitigate potential negative outcomes of the research, such as data and model documentation, data anonymization, responsible release, access control, and the reproducibility of findings? Yes, in Section 3.2
    9. (i)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? Yes
  2. 2.
    Additionally, if your study involves hypotheses testing…
    1. (a)
      Did you clearly state the assumptions underlying all theoretical results? NA
    2. (b)
      Have you provided justifications for all theoretical results? NA
    3. (c)
      Did you discuss competing hypotheses or theories that might challenge or complement your theoretical results? NA
    4. (d)
      Have you considered alternative mechanisms or explanations that might account for the same outcomes observed in your study? NA
    5. (e)
      Did you address potential biases or limitations in your theoretical framework? NA
    6. (f)
      Have you related your theoretical results to the existing literature in social science? NA
    7. (g)
      Did you discuss the implications of your theoretical results for policy, practice, or further research in the social science domain? NA
  3. 3.
    Additionally, if you are including theoretical proofs…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
  4. 4.
    Additionally, if you ran machine learning experiments…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? NA
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? NA
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? NA
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? NA
    5. (e)
      Do you justify how the proposed evaluation is sufficient and appropriate to the claims made? NA
    6. (f)
      Do you discuss what is “the cost“ of misclassification and fault (in)tolerance? NA
  5. 5.
    Additionally, if you are using existing assets (e.g., code, data, models) or curating/releasing new assets, without compromising anonymity
    1. (a)
      If your work uses existing assets, did you cite the creators? Yes, in Section 3.4
    2. (b)
      Did you mention the license of the assets? NA
    3. (c)
      Did you include any new assets in the supplemental material or as a URL? NA
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? Yes, in Section 3.2
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? Yes, in Section 3.2
    6. (f)
      If you are curating or releasing new datasets, did you discuss how you intend to make your datasets FAIR (see FORCE11 (2020))? NA
    7. (g)
      If you are curating or releasing new datasets, did you create a Datasheet for the Dataset (see Gebru et al. (2021))? NA
  6. 6.
    Additionally, if you used crowdsourcing or conducted research with human subjects, without compromising anonymity
    1. (a)
      Did you include the full text of instructions given to participants and screenshots? Yes, in Appendix B
    2. (b)
      Did you describe any potential participant risks, with mentions of Institutional Review Board (IRB) approvals? Yes, see Section 3.2 on Ethical Considerations
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? Yes, in Section 3.1
    4. (d)
      Did you discuss how data is stored, shared, and deidentified? Yes, in Section 3.2
Appendix B Questionnaires’ content
@@ B.1 Behavior questionnaire
    Do you think it is possible to revoke (cancel) previously granted mobile app access?
    • Yes
    • No
@@
    • User location – User’s contacts
    • SMS messages – MMS messages
    • User’s phone number – Email address
    • Username – Address
    • Camera – Audio
    • Phone call logs - – Storage
@@
Internet User’s Information Privacy Concerns based on a 7 Likert scale, see (Malhotra, Kim, and Agarwal 2004).
@@ Appendix C Codebooks
Table 1: Codebook for the thematic analysis performed on the answers to Q32 “How do you think location data could be misused by a malicious person?”
@@
Table 2: Codebook for the categorization performed on the answers to Q36 “What are the measures imposed by the GDPR that reduce these risks?”

### v3

@@ 3.5. Data analysis
The analysis was performed by two independent annotators, and a consensus was reached at the end of the process. For most fields, a full thematic analysis was neither required nor necessarily desirable; curation and normalisation of the data was however necessary. Indeed, stopping at the third step was enough for several fields, such as the reasons given to revoke location permission. A full thematic analysis was performed on the answers to Q32 “How do you think location data could be misused by a malicious person?”, the themes of which were loosely derived from (Citron and Solove, 2021) (Physical harms, Economic harms, Reputational harms, Informational integrity, Psychological harms), see Table 1 in Appendix B.
We regrouped our codes for Q36 “What are the measures imposed by the GDPR that reduce these risks?” under general legal categories (Rights, Consent, Clear purpose and transparency, Misc.) for statistical purposes, see Table 2 in Appendix B. The rationale behind Q36 was to evaluate participants’ knowledge about the GDPR.
3.6. General Statistics
@@
Note that our population is within average with respect to their privacy concerns ($\mu$ = 4.2, $\sigma$ = 1.05) [^12] See (Zufferey et al., 2023), for an example of less concerned participants ($\mu$ = 3.5, $\sigma$ = 1.6) and (Naeini et al., 2017) for more concerned participants with $\mu$ between 4.79 and 5.44 normalized to a [0-6] scale. in spite of young adults being generally less concerned about their privacy than older populations (as mentioned in Section 2).
Figure 4. Plot of the Internet User’s Information Privacy Concerns (IUIPC) distribution of participants.
4. Results
@@
4.1. Privacy risky practices
@@
Figure 5. Participants tend to underestimate the number of apps that have access to the location (Q3 against Q8).
@@
Figure 6. Participants tend not to think about disabling location access for apps that are not actively used or forget to do so (Q11).
@@
4.2. Participants are poorly aware of the risks
@@
Figure 7. Proportion of participants aware of source of data from which location can be inferred (Q12).
@@
Figure 8. Answers to Q13 “Do you think it is possible to infer the following information from the location?”.
@@
4.3. The risk demonstrator increased awareness of participants about privacy
@@
Figure 9. The most worrisome inferences are informational harms (i.e., surveillance, stalking, inference, data selling, targeted advertising) (Q32).
@@ 5.1. Recommendations
5.2. Limitations
First, our demographics do not represent a fair sample of digital users. Our participants were mostly males, all of them young adults educated in a top European engineering school, although with different backgrounds and technical skills. This bias in our representativeness played a part in the socio-economic backgrounds of participants, who are most likely to come from well-off social environments. Recruiting participants in universities often faces this limitation in academic studies on usages. For instance, (Kulyk et al., 2022) involved 40 participants mostly young and well-educated, while (Karegar, Pettersson, and Fischer-Hübner, 2020) involved 80 participants, mostly young, with 45% in the 18–25 age group and 20% between the ages of 26 and 35, a majority ($n=54$) were bachelor students or had their bachelor degrees. (Almuhimedi et al., 2015) involved 23 participants (65% female; ages 18–44, median=23), and (Kaushik et al., 2021) involved 392 participants ranging from 25 to 35.
@@ 5.2. Limitations
Second, the scope of our work must be restricted to a specific type of demonstrator, our results should not be generalized to all visualization and awareness tools. Moreover, the results largely presented self-reported and planned behaviors. Therefore, an interesting research avenue for future work is to study the actual behavioral changes introduced by the experiment of the tool, and long-term behavior analysis through longitudinal studies. Perceptions can also change according to the type of apps and their access to location data, as well as other sensitive personal data collected.
Third, participants used example data on the demonstration platform, which is perhaps less illustrative than if they had inspected their own location data. We envision working with personalized examples in the future. Our results also lack perspective on whether our participants were able to deeply understand the goal and the meaning of obfuscated traces and the epsilon number (see (Karegar, Alaqra, and Fischer-Hübner, 2022)).
6. Conclusion
@@
Acknowledgements
This work has been supported by the ANR 22-PECY-0002 IPOP (Interdisciplinary Project on Privacy) project of the Cybersecurity PEPR, and partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation. We would like to thank Farzaneh Karegar and Nataliia Bielova for their feedback on earlier drafts, and anonymous reviewers who helped us improve this article.
Appendix A Questionnaires’ content
@@ A.1. Behavior questionnaire
    Do you think it is possible to revoke (cancel) previously granted mobile app access? Yes/no
@@
    • User location
    • User’s contacts
    • SMS messages
    • MMS messages
    • User’s phone number
    • Email address
    • Username
    • Address
    • Camera
    • Audio
    • Phone call logs
    • Storage
@@
Internet User’s Information Privacy Concerns based on a 7 Likert scale, see [31].
@@ Appendix B Codebooks
Table 1. Codebook for the thematic analysis performed on the answers to Q32 “How do you think location data could be misused by a malicious person?”
@@
Table 2. Codebook for the categorization performed on the answers to Q36 “What are the measures imposed by the GDPR that reduce these risks?”